Can you give me the anser for x=2-2y
2x + 3y=4
Can you also give me a short explanation on how to do it. thanks.

Answer by checkley77(12844)

x=2-2y now substitute (2-2y) for x in the second equation & solve for y.
2x+3y=4
2(2-2y)+3y=4
4-4y+3y=4
-y=4-4
-y=0
y=0 ans.
x=2-2*0
x=2-0
x=2
Proof:
2*2+3*0=4
4+0=4
4=4